$SPAD/src/input setcmd.input

Timothy Daly December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)spool setcmd.output
)set message test on
)set message auto off
)clear all
--S 1 of 143
)set breakmode
--R-------------------------- The breakmode Option ---------------------------
--R
--R Description: execute break processing on error
--R
--R The breakmode option may be followed by any one of the following:
--R
--R    nobreak
--R    break
--R    query
--R -> resume
--R    fastlinks
--R    quit
--R
--R The current setting is indicated.
--R
--E 1
--S 2 of 143
)set compiler
--R                  Current Values of  compiler  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Routput       library in which to place compiled code    user.lib
--Rinput        controls libraries from which to load compiled code
--R
--E 2
--S 3 of 143
)set compiler
--R                  Current Values of  compiler  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Routput       library in which to place compiled code    user.lib
--Rinput        controls libraries from which to load compiled code
--R
--E 3

--S 4 of 143
)set compile input
--R---------------------------- The input Option -----------------------------
--R
--R Description: controls libraries from which to load compiled code
--R
--R )set compile input add library is used to tell AXIOM to add library to
--Rthe front of the path used to find compile code.
--R )set compile input drop library is used to tell AXIOM to remove library
--Rfrom this path.
--E 4
--S 5 of 143
)set compiler output
--R
--R---------------------------- The output Option ----------------------------
--R
--R Description: library in which to place compiled code
--R
--R )set compile output library is used to tell the compiler where to place
--Rcompiled code generated by the library compiler.  By default it goes
--Rin a file called user.lib in the current directory.
--E 5
--S 6 of 143 no longer supported
-- )set compiler args
--E 6
--S 7 of 143
)set expose
--R
--R---------------------------- The expose Option ----------------------------
--R
--R Description: control interpreter constructor exposure
--R
--R   The following groups are explicitly exposed in the current frame
--R      (called initial ):
--R                                   basic
--R                                categories
--R                                  naglink
--R                                   anna
--R
--R   The following constructors are explicitly exposed in the current
--R      frame:
--R               there are no explicitly exposed constructors
--R
--R   The following constructors are explicitly hidden in the current
--R      frame:
--R                there are no explicitly hidden constructors
--R

--E 7
--S 8 of 143
)set functions
--R                  Current Values of  functions  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rcache        number of function results to cache        0
--Rcompile      compile, don't just define function bodies on
--Rrecurrence   specially compile recurrence relations     on
--R
--E 8
--S 9 of 143
)set functions cache
--R
--R---------------------------- The cache Option -----------------------------
--R
--R Description: number of function results to cache
--R
--R )set functions cache is used to tell AXIOM how many
--R values computed by interpreter functions should be saved.  This
--R can save quite a bit of time in recursive functions, though one
--R must consider that the cached values will take up (perhaps
--R valuable) room in the workspace.
--R
--R The value given after cache must either be the word all or a positive integer.
--R This may be followed by any number of function names whose cache
--R sizes you wish to so set.  If no functions are given, the default
--R cache size is set.
--R
--R Examples:
--R   )set fun cache all         )set fun cache 10 f g Legendre
--R
--R In general, functions will cache no returned values.
--E 9
--S 10 of 143
)set functions compile
--R--------------------------- The compile Option ----------------------------
--R
--R Description: compile, don't just define function bodies
--R
--R The compile option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.

--R
--E 10
--S 11 of 143
)set functions recurrence
--R-------------------------- The recurrence Option --------------------------
--R
--R Description: specially compile recurrence relations
--R
--R The recurrence option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 11
--S 12 of 143
)set fortran
--R                   Current Values of  fortran  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rints2floats  where sensible, coerce integers to reals   on
--Rfortindent   the number of characters indented          6
--Rfortlength   the number of characters on a line         72
--Rtypedecs     print type and dimension lines             on
--Rdefaulttype  default generic type for FORTRAN object    REAL
--Rprecision    precision of generated FORTRAN objects     double
--Rintrinsic    whether to use INTRINSIC FORTRAN functions off
--Rexplength    character limit for FORTRAN expressions    1320
--Rsegment      split long FORTRAN expressions             on
--Roptlevel     FORTRAN optimisation level                 0
--Rstartindex   starting index for FORTRAN arrays          1
--Rcalling      options for external FORTRAN calls         ...
--R
--RVariables with current values of ... have further sub-options. For example,
--Rissue )set  calling to see what the options are for calling .
--RFor more information, issue )help set .
--E 12
--S 13 of 143
)set fortran ints2floats
--R------------------------- The ints2floats Option --------------------------
--R
--R Description: where sensible, coerce integers to reals
--R
--R The ints2floats option may be followed by any one of the following:
--R

--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 13
--S 14 of 143
)set fortran fortindent
--R-------------------------- The fortindent Option --------------------------
--R
--R Description: the number of characters indented
--R
--R The fortindent option may be followed by an integer in the range 0 to
--R  inclusive. The current setting is 6
--R
--E 14
--S 15 of 143
)set fortran fortlength
--R-------------------------- The fortlength Option --------------------------
--R
--R Description: the number of characters on a line
--R
--R The fortlength option may be followed by an integer in the range 1 to
--R  inclusive. The current setting is 72
--R
--E 15
--S 16 of 143
)set fortran typedecs
--R--------------------------- The typedecs Option ---------------------------
--R
--R Description: print type and dimension lines
--R
--R The typedecs option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 16
--S 17 of 143
)set fortran defaulttype
--R------------------------- The defaulttype Option --------------------------
--R
--R Description: default generic type for FORTRAN object
--R

--R The defaulttype option may be followed by any one of the following:
--R
--R -> REAL
--R    INTEGER
--R    COMPLEX
--R    LOGICAL
--R    CHARACTER
--R
--R The current setting is indicated.
--R
--E 17
--S 18 of 143
)set fortran precision
--R-------------------------- The precision Option ---------------------------
--R
--R Description: precision of generated FORTRAN objects
--R
--R The precision option may be followed by any one of the following:
--R
--R    single
--R -> double
--R
--R The current setting is indicated.
--R
--E 18
--S 19 of 143
)set fortran intrinsic
--R-------------------------- The intrinsic Option ---------------------------
--R
--R Description: whether to use INTRINSIC FORTRAN functions
--R
--R The intrinsic option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 19
--S 20 of 143
)set fortran explength
--R-------------------------- The explength Option ---------------------------
--R
--R Description: character limit for FORTRAN expressions
--R
--R The explength option may be followed by an integer in the range 0 to
--R  inclusive. The current setting is 1320

--S 21 of 143
)set fortran segment
--R--------------------------- The segment Option ----------------------------
--R
--R Description: split long FORTRAN expressions
--R
--R The segment option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 21
--S 22 of 143
)set fortran optlevel
--R--------------------------- The optlevel Option ---------------------------
--R
--R Description: FORTRAN optimisation level
--R
--R The optlevel option may be followed by an integer in the range 0 to
--R 2 inclusive. The current setting is 0
--R
--E 22
--S 23 of 143
)set fortran startindex
--R-------------------------- The startindex Option --------------------------
--R
--R Description: starting index for FORTRAN arrays
--R
--R The startindex option may be followed by an integer in the range 0 to
--R 1 inclusive. The current setting is 1
--R
--E 23
--S 24 of 143
)set fortran calling
--R                   Current Values of  calling  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rtempfile     set location of temporary data files       /tmp/
--Rdirectory    set location of generated FORTRAN files    ./
--Rlinker       linker arguments (e.g. libraries to search) -lxlf
--R

--E 24
--S 25 of 143
)set fortran calling tempfile
--R--------------------------- The tempfile Option ---------------------------
--R
--R Description: set location of temporary data files
--R
--R )set fortran calling tempfile  is used to tell AXIOM where
--R to place intermediate FORTRAN data files . This must be the
--R name of a valid existing directory to which you have permission
--R to write (including the final slash).
--R
--R Syntax:
--R   )set fortran calling tempfile DIRECTORYNAME
--R
--R The current setting is /tmp/
--E 25
--S 26 of 143
)set fortran calling directory
--R-------------------------- The directory Option ---------------------------
--R
--R Description: set location of generated FORTRAN files
--R
--R )set fortran calling directory  is used to tell AXIOM where
--R to place generated FORTRAN files. This must be the name
--R of a valid existing directory to which you have permission
--R to write (including the final slash).
--R
--R Syntax:
--R   )set fortran calling directory DIRECTORYNAME
--R
--R The current setting is ./
--E 26
--S 27 of 143
)set fortran calling linker
--R---------------------------- The linker Option ----------------------------
--R
--R Description: linker arguments (e.g. libraries to search)
--R
--R )set fortran calling linkerargs  is used to pass arguments to the linker
--R when using  mkFort  to create functions which call Fortran code.
--R For example, it might give a list of libraries to be searched,
--R and their locations.
--R The string is passed verbatim, so must be the correct syntax for
--R the particular linker being used.
--R
--R Example: )set fortran calling linker "-lxlf"

--R
--R The current setting is -lxlf
--E 27
--S 28 of 143 no longer supported
--)set kernel
--E 28
--S 29 of 143 no longer supported
--)set kernel warn
--E 29
--S 30 of 143 no longer supported
--)set kernel protect
--E 30
--S 31 of 143
)set hyperdoc
--R                  Current Values of  hyperdoc  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rfullscreen   use full screen for this facility          off
--Rmathwidth    screen width for history output            120
--R
--E 31
--S 32 of 143
)set hyperdoc fullscreen
--R-------------------------- The fullscreen Option --------------------------
--R
--R Description: use full screen for this facility
--R
--R The fullscreen option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 32
--S 33 of 143
)set hyperdoc mathwidth
--R-------------------------- The mathwidth Option ---------------------------
--R
--R Description: screen width for history output
--R
--R The mathwidth option may be followed by an integer in the range 0 to
--R  inclusive. The current setting is 120

--S 34 of 143
)set help
--R                    Current Values of  help  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rfullscreen   use fullscreen facility, if possible       on
--R
--E 34
--S 35 of 143
)set help fullscreen
--R-------------------------- The fullscreen Option --------------------------
--R
--R Description: use fullscreen facility, if possible
--R
--R The fullscreen option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 35
--S 36 of 143
)set history
--R--------------------------- The history Option ----------------------------
--R
--R Description: save workspace values in a history file
--R
--R The history option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 36
--S 37 of 143
)set messages
--R                  Current Values of  messages  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rany          print the internal type of objects of domain Any on

--Rautoload     print file auto-load messages              off
--Rbottomup     display bottom up modemap selection        off
--Rcoercion     display datatype coercion messages         off
--Rdropmap      display old map defn when replaced         off
--Rexpose       warning for unexposed functions            off
--Rfile         print msgs also to SPADMSG LISTING         off
--Rframe        display messages about frames              off
--Rhighlighting use highlighting in system messages        off
--Rinstant      present instantiation summary              off
--Rinsteach     present instantiation info                 off
--Rinterponly   say when function code is interpreted      on
--Rnaglink      show NAGLink messages                      on
--Rnumber       display message number with message        off
--Rprompt       set type of input prompt to display        step
--Rselection    display function selection msgs            off
--Rset          show )set setting after assignment         off
--Rstartup      display messages on start-up               on
--Rsummary      print statistics after computation         off
--Rtesting      print system testing header                on
--Rtime         print timings after computation            off
--Rtype         print type after computation               on
--Rvoid         print Void value when it occurs            off
--R
--E 37
--S 38 of 143
)set messages autoload
--R
--R--------------------------- The autoload Option ---------------------------
--R
--R Description: print file auto-load messages
--R
--R The autoload option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 38
--S 39 of 143
)set messages bottomup
--R--------------------------- The bottomup Option ---------------------------
--R
--R Description: display bottom up modemap selection
--R
--R The bottomup option may be followed by any one of the following:
--R
--R    on

--R -> off
--R
--R The current setting is indicated.
--R
--E 39
--S 40 of 143
)set messages coercion
--R--------------------------- The coercion Option ---------------------------
--R
--R Description: display datatype coercion messages
--R
--R The coercion option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 40
--S 41 of 143
)set messages dropmap
--R--------------------------- The dropmap Option ----------------------------
--R
--R Description: display old map defn when replaced
--R
--R The dropmap option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 41
--S 42 of 143
)set messages expose
--R---------------------------- The expose Option ----------------------------
--R
--R Description: warning for unexposed functions
--R
--R The expose option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 42

--S 43 of 143
)set messages file
--R----------------------------- The file Option -----------------------------
--R
--R Description: print msgs also to SPADMSG LISTING
--R
--R The file option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 43
--S 44 of 143
)set messages frame
--R---------------------------- The frame Option -----------------------------
--R
--R Description: display messages about frames
--R
--R The frame option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 44
--S 45 of 143
)set messages highlighting
--R------------------------- The highlighting Option -------------------------
--R
--R Description: use highlighting in system messages
--R
--R The highlighting option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 45
--S 46 of 143
)set messages instant
--R--------------------------- The instant Option ----------------------------
--R

--R Description: present instantiation summary
--R
--R The instant option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 46
--S 47 of 143
)set messages insteach
--R--------------------------- The insteach Option ---------------------------
--R
--R Description: present instantiation info
--R
--R The insteach option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 47
--S 48 of 143
)set messages interponly
--R-------------------------- The interponly Option --------------------------
--R
--R Description: say when function code is interpreted
--R
--R The interponly option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 48
--S 49 of 143
)set messages number
--R---------------------------- The number Option ----------------------------
--R
--R Description: display message number with message
--R
--R The number option may be followed by any one of the following:
--R
--R    on

--R -> off
--R
--R The current setting is indicated.
--R
--E 49
--S 50 of 143
)set messages prompt
--R---------------------------- The prompt Option ----------------------------
--R
--R Description: set type of input prompt to display
--R
--R The prompt option may be followed by any one of the following:
--R
--R    none
--R    frame
--R    plain
--R -> step
--R    verbose
--R
--R The current setting is indicated.
--R
--E 50
--S 51 of 143
)set messages selection
--R-------------------------- The selection Option ---------------------------
--R
--R Description: display function selection msgs
--R
--R The selection option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 51
--S 52 of 143
)set messages set
--R----------------------------- The set Option ------------------------------
--R
--R Description: show )set setting after assignment
--R
--R The set option may be followed by any one of the following:
--R
--R    on
--R -> off
--R

--R The current setting is indicated.
--R
--E 52
--S 53 of 143
)set messages startup
--R--------------------------- The startup Option ----------------------------
--R
--R Description: display messages on start-up
--R
--R The startup option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 53
--S 54 of 143
)set messages summary
--R--------------------------- The summary Option ----------------------------
--R
--R Description: print statistics after computation
--R
--R The summary option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 54
--S 55 of 143
)set messages testing
--R--------------------------- The testing Option ----------------------------
--R
--R Description: print system testing header
--R
--R The testing option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 55
--S 56 of 143

)set messages time
--R----------------------------- The time Option -----------------------------
--R
--R Description: print timings after computation
--R
--R The time option may be followed by any one of the following:
--R
--R    on
--R -> off
--R    long
--R
--R The current setting is indicated.
--R
--E 56
--S 57 of 143
)set messages type
--R----------------------------- The type Option -----------------------------
--R
--R Description: print type after computation
--R
--R The type option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 57
--S 58 of 143
)set messages void
--R----------------------------- The void Option -----------------------------
--R
--R Description: print Void value when it occurs
--R
--R The void option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 58
--S 59 of 143
)set messages any
--R----------------------------- The any Option ------------------------------
--R
--R Description: print the internal type of objects of domain Any

--R
--R The any option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 59
--S 60 of 143
)set messages naglink
--R--------------------------- The naglink Option ----------------------------
--R
--R Description: show NAGLink messages
--R
--R The naglink option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 60
--S 61 of 143
)set naglink host
--R----------------------------- The host Option -----------------------------
--R
--R Description: internet address of host for NAGLink
--R
--R )set naglink host is used to tell  AXIOM which  host to contact for
--R a NAGLink request. An Internet address should be supplied. The host
--R specified must be running the NAGLink daemon.
--R
--R The current setting is localhost
--E 61
--S 62 of 143
)set naglink persistence
--R------------------------- The persistence Option --------------------------
--R
--R Description: number of (fortran) functions to remember
--R
--R )set naglink persistence is used to tell  the  nagd  daemon how  many ASP
--R source and object files to keep around in case you reuse them. This helps
--R to avoid needless recompilations. The number specified should be a
--R non-negative integer.
--R
--R The current setting is 1

--E 62
--S 63 of 143
)set naglink messages
--R--------------------------- The messages Option ---------------------------
--R
--R Description: show NAGLink messages
--R
--R The messages option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 63
--S 64 of 143
)set naglink double
--R---------------------------- The double Option ----------------------------
--R
--R Description: enforce DOUBLE PRECISION ASPs
--R
--R The double option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 64
--S 65 of 143
)set output
--R
--R                   Current Values of  output  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rabbreviate   abbreviate type names                      off
--Ralgebra      display output in algebraic form           On:CONSOLE
--Rcharacters   choose special output character set        plain
--Rfortran      create output in FORTRAN format            Off:CONSOLE
--Rfraction     how fractions are formatted                vertical
--Rhtml         create output in HTML style                Off:CONSOLE
--Rlength       line length of output displays             77
--Rmathml       create output in MathML style              Off:CONSOLE
--Ropenmath     create output in OpenMath style            Off:CONSOLE
--Rscript       display output in SCRIPT formula format    Off:CONSOLE
--Rscripts      show subscripts,... linearly               off

--Rshoweditor   view output of )show in editor             off
--Rtex          create output in TeX style                 Off:CONSOLE
--R
--E 65
--S 66 of 143
)set output abbreviate
--R-------------------------- The abbreviate Option --------------------------
--R
--R Description: abbreviate type names
--R
--R The abbreviate option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 66
--S 67 of 143
)set output algebra
--R--------------------------- The algebra Option ----------------------------
--R
--R Description: display output in algebraic form
--R
--R )set output algebra is used to tell AXIOM to turn algebra-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output algebra <arg>
--R    where arg can be one of
--R  on          turn algebra printing on (default state)
--R  off         turn algebra printing off
--R  console     send algebra output to screen (default state)
--R  fp<.fe>     send algebra output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .spout.
--R
--RIf you wish to send the output to a file, you may need to issue this command
--Rtwice: once with on and once with the file name. For example, to send
--Ralgebra output to the file polymer.spout, issue the two commands
--R
--R  )set output algebra on
--R  )set output algebra polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  On:CONSOLE
--E 67

--S 68 of 143
)set output characters
--R-------------------------- The characters Option --------------------------
--R
--R Description: choose special output character set
--R
--R
--R The characters option may be followed by any one of the following:
--R
--R    default
--R -> plain
--R
--R The current setting is indicated within the list.  This option determines
--R the special characters used for algebraic output.  This is what the
--R current choice of special characters looks like:
--R   ulc is shown as +          urc is shown as +
--R   llc is shown as +          lrc is shown as +
--R   vbar is shown as |         hbar is shown as -
--R   quad is shown as ?         lbrk is shown as [
--R   rbrk is shown as ]         lbrc is shown as {
--R   rbrc is shown as }         ttee is shown as +
--R   btee is shown as +         rtee is shown as +
--R   ltee is shown as +         ctee is shown as +
--R   bslash is shown as \
--E 68
--S 69 of 143
)set output fortran
--R--------------------------- The fortran Option ----------------------------
--R
--R Description: create output in FORTRAN format
--R
--R )set output fortran is used to tell AXIOM to turn FORTRAN-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RAlso See: )set fortran
--R
--RSyntax:   )set output fortran <arg>
--R    where arg can be one of
--R  on          turn FORTRAN printing on
--R  off         turn FORTRAN printing off (default state)
--R  console     send FORTRAN output to screen (default state)
--R  fp<.fe>     send FORTRAN output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .sfort.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RFORTRAN output to the file polymer.sfort, issue the two commands
--R

--R  )set output fortran on
--R  )set output fortran polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 69
--S 70 of 143
)set output fraction
--R--------------------------- The fraction Option ---------------------------
--R
--R Description: how fractions are formatted
--R
--R The fraction option may be followed by any one of the following:
--R
--R -> vertical
--R    horizontal
--R
--R The current setting is indicated.
--R
--E 70
--S 71 of 143
)set output length
--R---------------------------- The length Option ----------------------------
--R
--R Description: line length of output displays
--R
--R The length option may be followed by an integer in the range 10 to
--R 245 inclusive. The current setting is 77
--R
--E 71
--S 72 of 143
)set output mathml
--R---------------------------- The mathml Option ----------------------------
--R
--R Description: create output in MathML style
--R
--R )set output mathml is used to tell AXIOM to turn MathML-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output mathml <arg>
--R    where arg can be one of
--R  on          turn MathML printing on
--R  off         turn MathML printing off (default state)
--R  console     send MathML output to screen (default state)
--R  fp<.fe>     send MathML output to file with file prefix fp and file

--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RMathML output to the file polymer.smml, issue the two commands
--R
--R  )set output mathml on
--R  )set output mathml polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 72
--S 73 of 143
)set output openmath
--R--------------------------- The openmath Option ---------------------------
--R
--R Description: create output in OpenMath style
--R
--R )set output openmath is used to tell AXIOM to turn OpenMath output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output openmath <arg>
--R    where arg can be one of
--R  on          turn OpenMath printing on
--R  off         turn OpenMath printing off (default state)
--R  console     send OpenMath output to screen (default state)
--R  fp<.fe>     send OpenMath output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .som.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--ROpenMath output to the file polymer.som, issue the two commands
--R
--R  )set output openmath on
--R  )set output openmath polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 73
--S 74 of 143
)set output script
--R---------------------------- The script Option ----------------------------
--R
--R Description: display output in SCRIPT formula format
--R

--R )set output script is used to tell AXIOM to turn IBM Script formula-style
--Routput printing on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output script <arg>
--R    where arg can be one of
--R  on          turn IBM Script formula printing on
--R  off         turn IBM Script formula printing off (default state)
--R  console     send IBM Script formula output to screen (default state)
--R  fp<.fe>     send IBM Script formula output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .sform.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RIBM Script formula output to the file polymer.sform, issue the two commands
--R
--R  )set output script on
--R  )set output script polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 74
--S 75 of 143
)set output scripts
--R--------------------------- The scripts Option ----------------------------
--R
--R Description: show subscripts,... linearly
--R
--R The scripts option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 75
--S 76 of 143
)set output showeditor
--R-------------------------- The showeditor Option --------------------------
--R
--R Description: view output of )show in editor
--R
--R The showeditor option may be followed by any one of the following:
--R
--R    on
--R -> off
--R

--S 77 of 143
)set output tex
--R----------------------------- The tex Option ------------------------------
--R
--R Description: create output in TeX style
--R
--R )set output tex is used to tell AXIOM to turn TeX-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output tex <arg>
--R    where arg can be one of
--R  on          turn TeX printing on
--R  off         turn TeX printing off (default state)
--R  console     send TeX output to screen (default state)
--R  fp<.fe>     send TeX output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RTeX output to the file polymer.stex, issue the two commands
--R
--R  )set output tex on
--R  )set output tex polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 77
--S 78 of 143
)set quit
--R----------------------------- The quit Option -----------------------------
--R
--R Description: protected or unprotected quit
--R
--R The quit option may be followed by any one of the following:
--R
--R -> protected
--R    unprotected
--R
--R The current setting is indicated.
--R
--E 78
--S 79 of 143

)set streams
--R                   Current Values of  streams  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rcalculate    specify number of elements to calculate    10
--Rshowall      display all stream elements computed       off
--R
--E 79
--S 80 of 143
)set streams calculate
--R-------------------------- The calculate Option ---------------------------
--R
--R Description: specify number of elements to calculate
--R
--R   )set streams calculate is used to tell AXIOM how many elements of a
--R      stream to calculate when a computation uses the stream. The value
--R      given after calculate must either be the word all or a positive
--R      integer.
--R
--R      The current setting is 10 .
--E 80
--S 81 of 143
)set streams showall
--R--------------------------- The showall Option ----------------------------
--R
--R Description: display all stream elements computed
--R
--R The showall option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 81
--S 82 of 143
)set system
--R                   Current Values of  system  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rfunctioncode show gen. LISP for functions when compiled off
--Roptimization show optimized LISP code                   off
--Rprettyprint  prettyprint BOOT func's as they compile    on
--R
--E 82

--S 83 of 143
)set system functioncode
--R------------------------- The functioncode Option -------------------------
--R
--R Description: show gen. LISP for functions when compiled
--R
--R The functioncode option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 83
--S 84 of 143
)set system optimization
--R------------------------- The optimization Option -------------------------
--R
--R Description: show optimized LISP code
--R
--R The optimization option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 84
--S 85 of 143
)set system prettyprint
--R------------------------- The prettyprint Option --------------------------
--R
--R Description: prettyprint BOOT func's as they compile
--R
--R The prettyprint option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 85
--S 86 of 143
)set userlevel
--R-------------------------- The userlevel Option ---------------------------
--R

--R Description: operation access level of system user
--R
--R The userlevel option may be followed by any one of the following:
--R
--R    interpreter
--R    compiler
--R -> development
--R
--R The current setting is indicated.
--R
--E 86
--S 87 of 143
)set output char
--R-------------------------- The characters Option --------------------------
--R
--R Description: choose special output character set
--R
--R
--R The characters option may be followed by any one of the following:
--R
--R    default
--R -> plain
--R
--R The current setting is indicated within the list.  This option determines
--R the special characters used for algebraic output.  This is what the
--R current choice of special characters looks like:
--R   ulc is shown as +          urc is shown as +
--R   llc is shown as +          lrc is shown as +
--R   vbar is shown as |         hbar is shown as -
--R   quad is shown as ?         lbrk is shown as [
--R   rbrk is shown as ]         lbrc is shown as {
--R   rbrc is shown as }         ttee is shown as +
--R   btee is shown as +         rtee is shown as +
--R   ltee is shown as +         ctee is shown as +
--R   bslash is shown as \
--E 87
--S 88 of 143
)set output char default
--E 88
--S 89 of 143
)set output char
--R The characters Option
--R
--R Description: choose special output character set
--R
--R
--R The characters option may be followed by any one of the following:

--R
--R -> default
--R    plain
--R
--R The current setting is indicated within the list.  This option determines
--R the special characters used for algebraic output.  This is what the
--R current choice of special characters looks like:
--R   ulc is shown as           urc is shown as
--R   llc is shown as           lrc is shown as
--R   vbar is shown as          hbar is shown as
--R   quad is shown as NIL       lbrk is shown as [
--R   rbrk is shown as ]         lbrc is shown as {
--R   rbrc is shown as }         ttee is shown as
--R   btee is shown as          rtee is shown as
--R   ltee is shown as          ctee is shown as
--R   bslash is shown as \
--E 89
--S 90 of 143
)set output char plain
--E 90
--S 91 of 143
)set output char
--R-------------------------- The characters Option --------------------------
--R
--R Description: choose special output character set
--R
--R
--R The characters option may be followed by any one of the following:
--R
--R    default
--R -> plain
--R
--R The current setting is indicated within the list.  This option determines
--R the special characters used for algebraic output.  This is what the
--R current choice of special characters looks like:
--R   ulc is shown as +          urc is shown as +
--R   llc is shown as +          lrc is shown as +
--R   vbar is shown as |         hbar is shown as -
--R   quad is shown as ?         lbrk is shown as [
--R   rbrk is shown as ]         lbrc is shown as {
--R   rbrc is shown as }         ttee is shown as +
--R   btee is shown as +         rtee is shown as +
--R   ltee is shown as +         ctee is shown as +
--R   bslash is shown as \
--E 91
--S 92 of 143
)set output fortran

--R--------------------------- The fortran Option ----------------------------
--R
--R Description: create output in FORTRAN format
--R
--R )set output fortran is used to tell AXIOM to turn FORTRAN-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RAlso See: )set fortran
--R
--RSyntax:   )set output fortran <arg>
--R    where arg can be one of
--R  on          turn FORTRAN printing on
--R  off         turn FORTRAN printing off (default state)
--R  console     send FORTRAN output to screen (default state)
--R  fp<.fe>     send FORTRAN output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .sfort.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RFORTRAN output to the file polymer.sfort, issue the two commands
--R
--R  )set output fortran on
--R  )set output fortran polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 92
--S 93 of 143
)set output fortran foo
--I   FORTRAN output will be written to file /research/test/foo.sfort .
--E 93
--S 94 of 143
)set output fortran
--R--------------------------- The fortran Option ----------------------------
--R
--R Description: create output in FORTRAN format
--R
--R )set output fortran is used to tell AXIOM to turn FORTRAN-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RAlso See: )set fortran
--R
--RSyntax:   )set output fortran <arg>
--R    where arg can be one of
--R  on          turn FORTRAN printing on

--R  off         turn FORTRAN printing off (default state)
--R  console     send FORTRAN output to screen (default state)
--R  fp<.fe>     send FORTRAN output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .sfort.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RFORTRAN output to the file polymer.sfort, issue the two commands
--R
--R  )set output fortran on
--R  )set output fortran polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  Off:/research/test/foo.sfort
--E 94
--S 95 of 143
)set output fortran on
--E 95
--S 96 of 143
)set output fortran
--R--------------------------- The fortran Option ----------------------------
--R
--R Description: create output in FORTRAN format
--R
--R )set output fortran is used to tell AXIOM to turn FORTRAN-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RAlso See: )set fortran
--R
--RSyntax:   )set output fortran <arg>
--R    where arg can be one of
--R  on          turn FORTRAN printing on
--R  off         turn FORTRAN printing off (default state)
--R  console     send FORTRAN output to screen (default state)
--R  fp<.fe>     send FORTRAN output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .sfort.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RFORTRAN output to the file polymer.sfort, issue the two commands
--R
--R  )set output fortran on
--R  )set output fortran polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.

--IThe current setting is:  On:/research/test/foo.sfort
--E 96
--S 97 of 143
)set output fortran append
--I   FORTRAN output will be written to file /research/test/NIL.sfort .
--E 97
--S 98 of 143
)set output fortran
--R--------------------------- The fortran Option ----------------------------
--R
--R Description: create output in FORTRAN format
--R
--R )set output fortran is used to tell AXIOM to turn FORTRAN-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RAlso See: )set fortran
--R
--RSyntax:   )set output fortran <arg>
--R    where arg can be one of
--R  on          turn FORTRAN printing on
--R  off         turn FORTRAN printing off (default state)
--R  console     send FORTRAN output to screen (default state)
--R  fp<.fe>     send FORTRAN output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .sfort.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RFORTRAN output to the file polymer.sfort, issue the two commands
--R
--R  )set output fortran on
--R  )set output fortran polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  On:/research/test/NIL.sfort
--E 98
--S 99 of 143
)set output fortran foo
--I   FORTRAN output will be written to file /research/test/foo.sfort .
--E 99
--S 100 of 143
)set output fortran
--R--------------------------- The fortran Option ----------------------------
--R
--R Description: create output in FORTRAN format

--R
--R )set output fortran is used to tell AXIOM to turn FORTRAN-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RAlso See: )set fortran
--R
--RSyntax:   )set output fortran <arg>
--R    where arg can be one of
--R  on          turn FORTRAN printing on
--R  off         turn FORTRAN printing off (default state)
--R  console     send FORTRAN output to screen (default state)
--R  fp<.fe>     send FORTRAN output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .sfort.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RFORTRAN output to the file polymer.sfort, issue the two commands
--R
--R  )set output fortran on
--R  )set output fortran polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  On:/research/test/foo.sfort
--E 100
--S 101 of 143
)set output algebra
--R--------------------------- The algebra Option ----------------------------
--R
--R Description: display output in algebraic form
--R
--R )set output algebra is used to tell AXIOM to turn algebra-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output algebra <arg>
--R    where arg can be one of
--R  on          turn algebra printing on (default state)
--R  off         turn algebra printing off
--R  console     send algebra output to screen (default state)
--R  fp<.fe>     send algebra output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .spout.
--R
--RIf you wish to send the output to a file, you may need to issue this command
--Rtwice: once with on and once with the file name. For example, to send
--Ralgebra output to the file polymer.spout, issue the two commands
--R
--R  )set output algebra on

--R  )set output algebra polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  On:CONSOLE
--E 101
--S 102 of 143
)set output algebra foo
--E 102
--S 103 of 143
)set output algebra
--R--------------------------- The algebra Option ----------------------------
--R
--R Description: display output in algebraic form
--R
--R )set output algebra is used to tell AXIOM to turn algebra-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output algebra <arg>
--R    where arg can be one of
--R  on          turn algebra printing on (default state)
--R  off         turn algebra printing off
--R  console     send algebra output to screen (default state)
--R  fp<.fe>     send algebra output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .spout.
--R
--RIf you wish to send the output to a file, you may need to issue this command
--Rtwice: once with on and once with the file name. For example, to send
--Ralgebra output to the file polymer.spout, issue the two commands
--R
--R  )set output algebra on
--R  )set output algebra polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  On:/tmp/foo.spout
--E 103
--S 104 of 143
)set output algebra off
--E 104
--S 105 of 143
)set output algebra
--R--------------------------- The algebra Option ----------------------------
--R
--R Description: display output in algebraic form

--R
--R )set output algebra is used to tell AXIOM to turn algebra-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output algebra <arg>
--R    where arg can be one of
--R  on          turn algebra printing on (default state)
--R  off         turn algebra printing off
--R  console     send algebra output to screen (default state)
--R  fp<.fe>     send algebra output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .spout.
--R
--RIf you wish to send the output to a file, you may need to issue this command
--Rtwice: once with on and once with the file name. For example, to send
--Ralgebra output to the file polymer.spout, issue the two commands
--R
--R  )set output algebra on
--R  )set output algebra polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  Off:/tmp/foo.spout
--E 105
--S 106 of 143
)set output algebra console
--E 106
--S 107 of 143
)set output algebra
--R--------------------------- The algebra Option ----------------------------
--R
--R Description: display output in algebraic form
--R
--R )set output algebra is used to tell AXIOM to turn algebra-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output algebra <arg>
--R    where arg can be one of
--R  on          turn algebra printing on (default state)
--R  off         turn algebra printing off
--R  console     send algebra output to screen (default state)
--R  fp<.fe>     send algebra output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .spout.
--R
--RIf you wish to send the output to a file, you may need to issue this command
--Rtwice: once with on and once with the file name. For example, to send
--Ralgebra output to the file polymer.spout, issue the two commands

--R
--R  )set output algebra on
--R  )set output algebra polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 107
--S 108 of 143
)set output algebra on
--E 108
--S 109 of 143
)set output algebra
--R--------------------------- The algebra Option ----------------------------
--R
--R Description: display output in algebraic form
--R
--R )set output algebra is used to tell AXIOM to turn algebra-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output algebra <arg>
--R    where arg can be one of
--R  on          turn algebra printing on (default state)
--R  off         turn algebra printing off
--R  console     send algebra output to screen (default state)
--R  fp<.fe>     send algebra output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .spout.
--R
--RIf you wish to send the output to a file, you may need to issue this command
--Rtwice: once with on and once with the file name. For example, to send
--Ralgebra output to the file polymer.spout, issue the two commands
--R
--R  )set output algebra on
--R  )set output algebra polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  On:CONSOLE
--E 109
--S 110 of 143
)set output mathml
--R---------------------------- The mathml Option ----------------------------
--R
--R Description: create output in MathML style
--R
--R )set output mathml is used to tell AXIOM to turn MathML-style output

--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output mathml <arg>
--R    where arg can be one of
--R  on          turn MathML printing on
--R  off         turn MathML printing off (default state)
--R  console     send MathML output to screen (default state)
--R  fp<.fe>     send MathML output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RMathML output to the file polymer.smml, issue the two commands
--R
--R  )set output mathml on
--R  )set output mathml polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 110
--S 111 of 143
)set output mathml foo
--I   MathML output will be written to file /research/test/foo.smml .
--E 111
--S 112 of 143
)set output mathml
--R---------------------------- The mathml Option ----------------------------
--R
--R Description: create output in MathML style
--R
--R )set output mathml is used to tell AXIOM to turn MathML-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output mathml <arg>
--R    where arg can be one of
--R  on          turn MathML printing on
--R  off         turn MathML printing off (default state)
--R  console     send MathML output to screen (default state)
--R  fp<.fe>     send MathML output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RMathML output to the file polymer.smml, issue the two commands
--R

--R  )set output mathml on
--R  )set output mathml polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  Off:/research/test/foo.smml
--E 112
--S 113 of 143
)set output mathml on
--E 113
--S 114 of 143
)set output mathml
--R---------------------------- The mathml Option ----------------------------
--R
--R Description: create output in MathML style
--R
--R )set output mathml is used to tell AXIOM to turn MathML-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output mathml <arg>
--R    where arg can be one of
--R  on          turn MathML printing on
--R  off         turn MathML printing off (default state)
--R  console     send MathML output to screen (default state)
--R  fp<.fe>     send MathML output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RMathML output to the file polymer.smml, issue the two commands
--R
--R  )set output mathml on
--R  )set output mathml polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  On:/research/test/foo.smml
--E 114
--S 115 of 143
)set output mathml console
--E 115
--S 116 of 143
)set output mathml
--R---------------------------- The mathml Option ----------------------------
--R

--R Description: create output in MathML style
--R
--R )set output mathml is used to tell AXIOM to turn MathML-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output mathml <arg>
--R    where arg can be one of
--R  on          turn MathML printing on
--R  off         turn MathML printing off (default state)
--R  console     send MathML output to screen (default state)
--R  fp<.fe>     send MathML output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RMathML output to the file polymer.smml, issue the two commands
--R
--R  )set output mathml on
--R  )set output mathml polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  On:CONSOLE
--E 116
--S 117 of 143
)set output openmath
--R--------------------------- The openmath Option ---------------------------
--R
--R Description: create output in OpenMath style
--R
--R )set output openmath is used to tell AXIOM to turn OpenMath output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output openmath <arg>
--R    where arg can be one of
--R  on          turn OpenMath printing on
--R  off         turn OpenMath printing off (default state)
--R  console     send OpenMath output to screen (default state)
--R  fp<.fe>     send OpenMath output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .som.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--ROpenMath output to the file polymer.som, issue the two commands
--R
--R  )set output openmath on
--R  )set output openmath polymer

--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 117
--S 118 of 143
)set output openmath on
--E 118
--S 119 of 143
)set output openmath
--R--------------------------- The openmath Option ---------------------------
--R
--R Description: create output in OpenMath style
--R
--R )set output openmath is used to tell AXIOM to turn OpenMath output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output openmath <arg>
--R    where arg can be one of
--R  on          turn OpenMath printing on
--R  off         turn OpenMath printing off (default state)
--R  console     send OpenMath output to screen (default state)
--R  fp<.fe>     send OpenMath output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .som.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--ROpenMath output to the file polymer.som, issue the two commands
--R
--R  )set output openmath on
--R  )set output openmath polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  On:CONSOLE
--E 119
--S 120 of 143
)set output openmath foo
--I   OpenMath output will be written to file /research/test/foo.som .
--E 120
--S 121 of 143
)set output openmath
--R--------------------------- The openmath Option ---------------------------
--R
--R Description: create output in OpenMath style

--R
--R )set output openmath is used to tell AXIOM to turn OpenMath output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output openmath <arg>
--R    where arg can be one of
--R  on          turn OpenMath printing on
--R  off         turn OpenMath printing off (default state)
--R  console     send OpenMath output to screen (default state)
--R  fp<.fe>     send OpenMath output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .som.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--ROpenMath output to the file polymer.som, issue the two commands
--R
--R  )set output openmath on
--R  )set output openmath polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  On:/research/test/foo.som
--E 121
--S 122 of 143
)set output openmath off
--E 122
--S 123 of 143
)set output openmath
--R--------------------------- The openmath Option ---------------------------
--R
--R Description: create output in OpenMath style
--R
--R )set output openmath is used to tell AXIOM to turn OpenMath output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output openmath <arg>
--R    where arg can be one of
--R  on          turn OpenMath printing on
--R  off         turn OpenMath printing off (default state)
--R  console     send OpenMath output to screen (default state)
--R  fp<.fe>     send OpenMath output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .som.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--ROpenMath output to the file polymer.som, issue the two commands

--R
--R  )set output openmath on
--R  )set output openmath polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  Off:/research/test/foo.som
--E 123
--S 124 of 143
)set output openmath console
--E 124
--S 125 of 143
)set output openmath
--R--------------------------- The openmath Option ---------------------------
--R
--R Description: create output in OpenMath style
--R
--R )set output openmath is used to tell AXIOM to turn OpenMath output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output openmath <arg>
--R    where arg can be one of
--R  on          turn OpenMath printing on
--R  off         turn OpenMath printing off (default state)
--R  console     send OpenMath output to screen (default state)
--R  fp<.fe>     send OpenMath output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .som.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--ROpenMath output to the file polymer.som, issue the two commands
--R
--R  )set output openmath on
--R  )set output openmath polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 125
--S 126 of 143
)set output script
--R---------------------------- The script Option ----------------------------
--R
--R Description: display output in SCRIPT formula format
--R
--R )set output script is used to tell AXIOM to turn IBM Script formula-style

--Routput printing on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output script <arg>
--R    where arg can be one of
--R  on          turn IBM Script formula printing on
--R  off         turn IBM Script formula printing off (default state)
--R  console     send IBM Script formula output to screen (default state)
--R  fp<.fe>     send IBM Script formula output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .sform.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RIBM Script formula output to the file polymer.sform, issue the two commands
--R
--R  )set output script on
--R  )set output script polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 126
--S 127 of 143
)set output script on
--E 127
--S 128 of 143
)set output script
--R---------------------------- The script Option ----------------------------
--R
--R Description: display output in SCRIPT formula format
--R
--R )set output script is used to tell AXIOM to turn IBM Script formula-style
--Routput printing on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output script <arg>
--R    where arg can be one of
--R  on          turn IBM Script formula printing on
--R  off         turn IBM Script formula printing off (default state)
--R  console     send IBM Script formula output to screen (default state)
--R  fp<.fe>     send IBM Script formula output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .sform.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RIBM Script formula output to the file polymer.sform, issue the two commands
--R
--R  )set output script on

--R  )set output script polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  On:CONSOLE
--E 128
--S 129 of 143
)set output script foo
--I   IBM Script formula output will be written to file
--I      /research/test/foo.sform .
--E 129
--S 130 of 143
)set output script
--R---------------------------- The script Option ----------------------------
--R
--R Description: display output in SCRIPT formula format
--R
--R )set output script is used to tell AXIOM to turn IBM Script formula-style
--Routput printing on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output script <arg>
--R    where arg can be one of
--R  on          turn IBM Script formula printing on
--R  off         turn IBM Script formula printing off (default state)
--R  console     send IBM Script formula output to screen (default state)
--R  fp<.fe>     send IBM Script formula output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .sform.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RIBM Script formula output to the file polymer.sform, issue the two commands
--R
--R  )set output script on
--R  )set output script polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  On:/research/test/foo.sform
--E 130
--S 131 of 143
)set output script console
--E 131
--S 132 of 143
)set output script
--R---------------------------- The script Option ----------------------------

--R
--R Description: display output in SCRIPT formula format
--R
--R )set output script is used to tell AXIOM to turn IBM Script formula-style
--Routput printing on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output script <arg>
--R    where arg can be one of
--R  on          turn IBM Script formula printing on
--R  off         turn IBM Script formula printing off (default state)
--R  console     send IBM Script formula output to screen (default state)
--R  fp<.fe>     send IBM Script formula output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .sform.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RIBM Script formula output to the file polymer.sform, issue the two commands
--R
--R  )set output script on
--R  )set output script polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  On:CONSOLE
--E 132
--S 133 of 143
)set output script off
--E 133
--S 134 of 143
)set output script
--R---------------------------- The script Option ----------------------------
--R
--R Description: display output in SCRIPT formula format
--R
--R )set output script is used to tell AXIOM to turn IBM Script formula-style
--Routput printing on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output script <arg>
--R    where arg can be one of
--R  on          turn IBM Script formula printing on
--R  off         turn IBM Script formula printing off (default state)
--R  console     send IBM Script formula output to screen (default state)
--R  fp<.fe>     send IBM Script formula output to file with file prefix fp
--R              and file extension .fe. If not given, .fe defaults to .sform.
--R
--RIf you wish to send the output to a file, you must issue this command

--Rtwice: once with on and once with the file name. For example, to send
--RIBM Script formula output to the file polymer.sform, issue the two commands
--R
--R  )set output script on
--R  )set output script polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 134
--S 135 of 143
)set output tex
--R----------------------------- The tex Option ------------------------------
--R
--R Description: create output in TeX style
--R
--R )set output tex is used to tell AXIOM to turn TeX-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output tex <arg>
--R    where arg can be one of
--R  on          turn TeX printing on
--R  off         turn TeX printing off (default state)
--R  console     send TeX output to screen (default state)
--R  fp<.fe>     send TeX output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RTeX output to the file polymer.stex, issue the two commands
--R
--R  )set output tex on
--R  )set output tex polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 135
--S 136 of 143
)set output tex on
--E 136
--S 137 of 143
)set output tex
--R----------------------------- The tex Option ------------------------------
--R
--R Description: create output in TeX style

--R
--R )set output tex is used to tell AXIOM to turn TeX-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output tex <arg>
--R    where arg can be one of
--R  on          turn TeX printing on
--R  off         turn TeX printing off (default state)
--R  console     send TeX output to screen (default state)
--R  fp<.fe>     send TeX output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RTeX output to the file polymer.stex, issue the two commands
--R
--R  )set output tex on
--R  )set output tex polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  On:CONSOLE
--E 137
--S 138 of 143
)set output tex foo
--I   TeX output will be written to file /research/test/foo.stex .
--E 138
--S 139 of 143
)set output tex
--R----------------------------- The tex Option ------------------------------
--R
--R Description: create output in TeX style
--R
--R )set output tex is used to tell AXIOM to turn TeX-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output tex <arg>
--R    where arg can be one of
--R  on          turn TeX printing on
--R  off         turn TeX printing off (default state)
--R  console     send TeX output to screen (default state)
--R  fp<.fe>     send TeX output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send

--RTeX output to the file polymer.stex, issue the two commands
--R
--R  )set output tex on
--R  )set output tex polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  On:/research/test/foo.stex
--E 139
--S 140 of 143
)set output tex off
--E 140
--S 141 of 143
)set output tex
--R----------------------------- The tex Option ------------------------------
--R
--R Description: create output in TeX style
--R
--R )set output tex is used to tell AXIOM to turn TeX-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output tex <arg>
--R    where arg can be one of
--R  on          turn TeX printing on
--R  off         turn TeX printing off (default state)
--R  console     send TeX output to screen (default state)
--R  fp<.fe>     send TeX output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RTeX output to the file polymer.stex, issue the two commands
--R
--R  )set output tex on
--R  )set output tex polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--IThe current setting is:  Off:/research/test/foo.stex
--E 141
--S 142 of 143
)set output tex console
--E 142
--S 143 of 143
)set output tex

--R----------------------------- The tex Option ------------------------------
--R
--R Description: create output in TeX style
--R
--R )set output tex is used to tell AXIOM to turn TeX-style output
--Rprinting on and off, and where to place the output.  By default, the
--Rdestination for the output is the screen but printing is turned off.
--R
--RSyntax:   )set output tex <arg>
--R    where arg can be one of
--R  on          turn TeX printing on
--R  off         turn TeX printing off (default state)
--R  console     send TeX output to screen (default state)
--R  fp<.fe>     send TeX output to file with file prefix fp and file
--R              extension .fe. If not given, .fe defaults to .stex.
--R
--RIf you wish to send the output to a file, you must issue this command
--Rtwice: once with on and once with the file name. For example, to send
--RTeX output to the file polymer.stex, issue the two commands
--R
--R  )set output tex on
--R  )set output tex polymer
--R
--RThe output is placed in the directory from which you invoked AXIOM or
--Rthe one you set with the )cd system command.
--RThe current setting is:  Off:CONSOLE
--E 143
)spool
)lisp (bye)

———-

References^{[1] nothing}

52